A roller coaster begins at 7 m/s at the top of a 23 m hill. Find the velocity at a point 10 m above the bottom of the hill. Hint: (KE+PE) at point 1 = (KE+PE) at point 2
Answer #1

Let, intial position be denoted by 1 and final by 2
use,
K.E1 + P.E1 = K.E2 + P.E2
(0.5*m*v1^2) + m*g*h1 = (0.5*m*v2^2) + m*g*h2
(0.5*v1^2) + g*h1 = (0.5*v2^2) + g*h2
(0.5*7^2) + 9.8*23 = (0.5*v2^2) + 9.8*10
24.5 + 225.4 = (0.5*v2^2) + 98
151.9 = 0.5*v2^2
v2^2 = 303.8
v2 = 17.4 m/s

velocity at point 10 m above the bottom = 17.4 m/s
Answer: 17.4 m/s